OpenStudy (anonymous):

Simplfy by taking roots of the numerator and the demoniator. 3(sqrt)1000x^5/ y^3

OpenStudy (saifoo.khan):

Like that? \[\frac{3 \sqrt{1000x^5}}{y^3}\]

OpenStudy (anonymous):

Thats how it should look but now I need to know how to simplfy and take the roots of the numerator and demoniator.

OpenStudy (saifoo.khan):

\[\frac{30 x^2 \sqrt{10x}}{y^3}\]

OpenStudy (anonymous):

Could you by chance help me with this problem? 9b^2/5b^2-20b+20 * 5b-10/3b

OpenStudy (saifoo.khan):

Sure! :D

OpenStudy (saifoo.khan):

\[\frac{9b^2}{5b^2-20b+20} \times \frac{ 5b-10}{3b}\]Just correct me is it like that?

OpenStudy (anonymous):

yes exactly :)

OpenStudy (saifoo.khan):

Let me solve... :) If im not wrong you are doing mathlab..

OpenStudy (saifoo.khan):

Dang! i copied incorrect. just a sec.

OpenStudy (saifoo.khan):

\[\frac{3b}{b-2}\]

OpenStudy (anonymous):

thanks a million!

OpenStudy (saifoo.khan):

You welcome. Do you have anymore?